OpenStudy (anonymous):

(2x-2)/4x is taht simplified all the way

OpenStudy (anonymous):

take 2 common from numerator

OpenStudy (anonymous):

You can factor 2 from the numerator..

OpenStudy (anonymous):

now cancel 2 in numerator with 4 in denominator

OpenStudy (anonymous):

\[\frac{2x - 2}{4x} \implies \frac{2(x-1)}{4x}\]

OpenStudy (anonymous):

(x-1)/2x??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well Done..

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

u guys roock

OpenStudy (anonymous):

Welcome dear..

OpenStudy (unklerhaukus):

keep going\[(x-1)/2x\Rightarrow x/2x-1/2x=\]

OpenStudy (anonymous):

If you want to then you can proceed further also As Uncle Rocks is saying..